Find the particular solution to the equation

y''+25y=sec(5x)

You may omit absolute value signs.
yp(x)=
Answer #1

Given DE is yll +255 sec (50). The given differential equation is (2²+5²) y = sec (50) where D = dn The auxiliary equation isdy an -ay =d This is linases equation whose interording factor is an c -an ye Seconda ean Speron din Q = ean Beando proved D-Replacing é by ad we get 1 secsh -15h е ग Dats (6) log (cossa) Therefore particulan solution I secsn ܂ ܐܡ joi Dois secsn - 1